opposite 19
hypotenuses x
adjacent 17
adjacent
find the value of x

Answer by engsaoudah(9)

Using Pythagoras theorem
adjacent^2+opposite^2=hypotenuse^2
opposite=19
adjacent=17
17^2+ 19^2=hyp^2
hyp2=650
hyp= √650
hypotenuse=25.495
hypotenuse=25.5(1d.P)